OpenStudy (anonymous):

Simplify √-175 using the imaginary number i.

OpenStudy (saifoo.khan):

what's \(\sqrt{-1} \)?

OpenStudy (anonymous):

sqrt(ab)= sqrt(a)*sqrt(b) therefore with that being said sqrt(-1)*sqrt(175) = isqrt(175) because sqrt(-1) = i

OpenStudy (anonymous):

and because we know 175 = 25*7 so we can simplify more and write isqrt(25)sqrt(7) = 5isqrt(7)